OpenStudy (anonymous):

A painter leans a 20-ft ladder against a building. The base of the ladder is 12 ft from the building. To the nearest foot, how high on the building does the ladder reach?

OpenStudy (anonymous):

just to be sure, is the illustration like this? 

OpenStudy (anonymous):

it has no picture, and im not sure how it would look like :/ @Hidden_Twilight

OpenStudy (anonymous):

well, if it's like that. just use the Pythagorean theorem which is b = sqrt(c^2 + a^2).

OpenStudy (anonymous):

if only i kknew how to do that :/

OpenStudy (anonymous):



OpenStudy (anonymous):

let a = 12 ft and c = 20 ft were looking for b therefore: b^2 = c^2 - a^2 try to solve?

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

A painter leans a 20-ft ladder against a building. The base of the ladder is 12 ft from the building. To the nearest foot, how high on the building does the ladder reach

OpenStudy (anonymous):

b^2=c^2-a^2 b^2=20^2-12^2 b^2=400-144 b^2=256 b=sq rt of 256 =16ft Therefore the ladder will reach 16ft on the building